A solid cylindrical conducting shell of inner radius a -4 cm and outer radius b 6.7 cm has its axis aligned with the z axis as shown. It carries a uniformly distributed current l2- 7.8 A in the positive z-direction. An inifinte conducting wire is located along the z-axis and carries a current i1 = 3.6 A in the negative z-direction. , R(7074, 707d) R(7074.7074) Pido) 1) What is By(P), the y-component of the magnetic field at point P, located a distance d- 22 cm from the origin along the x-axis as shown? T Submit
Answer #1

1 )

B = \mu0 i / 2 \pi d

i = i2 - i1

i = 7.8 - 3.6

i = 4.2 A

By(P) = 4 X 3.14 X 10-7 X 4.2 / 2 X 3.14 X 0.21

By(P) = 4 X 10-6 T

2 )

\tiny \int_{P}^{S}B .d L = B X 3.14 X 0.21 / 4

\tiny \int_{P}^{S}B .d L = 4 X 10-6 X 3.14 X 0.21 / 4

\tiny \int_{P}^{S}B .d L = 6.59 X 10-7 T . m

3 )

i2,enc = i2 ( 3.14 ( r2 - a2 ) / 3.14 ( b2 - a2 ) )

i2,enc = 7.8 X ( ( 4.62 - 42 ) / ( 6.72 - 42 ))

i2,enc = 1.39

ienc = i1 + i2,enc

ienc = -3.6 + 1.39

ienc = -2.21 A

By(T) = - 4 X 3.14 X 10-7 X ( -2.21) / 2 X 3.14 X 0.046 )

By(T) = 9.6 X 10-6 T

4 )

\tiny \int_{S}^{P}B .d L = - 6.59 X 10-7 T . m

( same way limits are changed )

5 )

B ( 2 cm , 0 ) remains the same